OpenStudy (anonymous):

Find an equation of the line passing through the given points. (4,-8) and (-3,3) F(x)=

OpenStudy (anonymous):

Can someone please show me how to do this problem

OpenStudy (anonymous):

I am stressing out over this and cant figure it out

OpenStudy (anonymous):

use two point form for this

OpenStudy (callisto):

\[\frac{y-y_1}{x-x_1} = \frac{y_2-y_1}{x_2-x_1}\]Put \((x_1, y_1)\) =(4,-8) and \((x_2, y_2)\) = (-3,3) into above formula.\[\frac{y-(-8)}{x-4} = \frac{3-(-8)}{(-3)-4}\]\[\frac{y+8}{x-4} = \frac{11}{-7}\]\[y+8 = \frac{11}{-7}(x-4)\]\[y = \frac{11}{-7}(x-4)-8\] Simplify it. Change y to f(x) .